Help plz dont understand how to do it

Review Conceptual Example 2 as background for thisproblem.One kind of slingshot consists of a pocket thatholds a pebble and is whirled on a circle of radiusr.Thepebble is released from the circle at the angle θ, so that itwill hit the target.The distance to the target from thecenter of the circle isd.The circular path is parellel tothe ground, and the target lies in the plane of the circle. The distance d is ten times radius r.Ignore theeffect ofgravity in pulling the stone downward after it is released and findthe angle θ.
Answer #1
Given r is the radius and d is the distance to the target fromthe center of the circle.And given d = 10 rFrom figure cosα = ( r / d )=( r / 10r )=(1 / 10 )=(0.1 )α= 84.20Form the example the angle in the left corner below is350Therefore α + θ + 350 =1800θ = 1800- 350 - 84.20=60.80